OpenStudy (anonymous):

HELP PLEASE! For the function y = x(x-2)(x + 2), its y-intercept is y = its x-intercepts are x= , Note: If there is more than one answer enter them separated by commas. If there are none, enter none . When x\to \infty, y \to \infty (Input + or - for the answer) When x\to -\infty, y \to \infty (Input + or - for the answer)

OpenStudy (xapproachesinfinity):

y-intercept is when x=0 so y=0 x intercepts y=0 so x(x-2)(x+2)=0 x=0, x=2, x=-2 are xintercepts

OpenStudy (xapproachesinfinity):

\[\lim_{x\to \infty}y=\infty \] \[\lim _{x\to -\infty }y= -\infty\]